OpenStudy (anonymous):

2xy+1=xy+3;x=2 help Please!!

OpenStudy (bahrom7893):

x=2 2*2y+1=2y+3 4y+1=2y+3 2y=2 y=1

OpenStudy (theviper):

Any doubt now @jazzermains ??

OpenStudy (anonymous):

Nah